OpenStudy (logan13):

What is the exact volume of a sphere with a diameter of 36 in.?

OpenStudy (logan13):

http://prntscr.com/adnz4s

OpenStudy (logan13):

so its 904.78

OpenStudy (anonymous):

no

OpenStudy (anonymous):

V≈24429.02

OpenStudy (anonymous):

that was a mistake.

OpenStudy (anonymous):

the real answer is V≈24429.02

OpenStudy (logan13):

@ganeshie8

OpenStudy (pucusana):

can any of you help me with a question

OpenStudy (pucusana):

please?

jhonyy9 (jhonyy9):

than the diameter is 36 in so from 36/2 = 18 in resulte the length of radius formula for volume of a sphere = (4pir^3/3 V = (4*3,14*18^3)/3 = ? hope this will help you sure easy to get the right answer